FHC...a “Modest Proposal”

I would like to laud Professor Thurow for pointing out the underlying theme behind the Freshman Housing Committee Proposal. During Wednesday’s faculty meeting, he clarified this point with the following statement:

(get quote from ted)

The FHC proposal certainly goes a long way toward moving these decisions out of the incapable hands of childish freshmen and into the more experienced hand of the Institute. After living in carefully selected institute housing for a year, undergraduates will undoubtedly be better equipped to make intelligent, adult-like decisions about where to live [alternately: ...to make living group decisions more in line with mainstream institute doctrine].

Perhaps the next issue the Institute should consider is that of academic major selection. It is unlikely that students incapable of making adult decisions about their own living accomodations will mature sufficiently in one year at MIT to make adult decisions about more important issues such as that of chosing an academic major. If, instead, the Institute would randomly assign each student to a major at the end of his or her freshman year, this problem would be allileviated in the same manner as the housing decision. Additionally, this would offer the institute and students a number of advantages:

• Eliminate over/undercrowding of Departments – the Institute will be able to determine and control the number of undergraduate majors in each department; this will make it easier to allocate resources among the departments.

• Promote Departmental Diversity – Each department will be guaranteed a random mix of students. No longer will Course Six have to suffer with a nerdly stereotype because the Nerds will be evenly distributed among all departments.

• Better Serve Industrial and Government Sponsors – Since the Institute could then tightly control the number of graduates per year in each discipline, it can more closely match the needs of its various sponsors.

This, of course, would just be the beginning. Once these changes are implemented the institute could continue its effort to protect the undergraduates from their own inability to make adult choices in other areas. For example, the Institute could randomly preassign every student a boy/girl-friend. This would allieviate undergraduates from the stress of mate selection allowing them to concentrate more fully on their studies. To further promote diversity, the Institute could randomly pre-assign 10% of the students to a partner of the same sex.

Clearly, the implementation of the Freshman Housing Committee Proposal would usher in a new era at MIT, and things would never be the same again.